In a city 1/6 of the people were 60 years old 1/4 were under 18 years if the total number of people in the city is 6000 then how many people were between 18 and 60years
Answer by rfer(16322)

6000/6=1000 60 yrs
6000/4=1500 under 18
6000-1000-1500=3500 between 18 & 60